OpenStudy (anonymous):

I am really not sure how to simplify this: 2x(3x+4)^4/3+4x^2(3x+4)^1/3?

OpenStudy (anonymous):

Are they raised to fractional powers?

OpenStudy (anonymous):

For example,\[2x(3x+4)^{4/3}\]

OpenStudy (anonymous):

they are raised to fractional powers....I know there has to be someway the (3x+4) comes out. I'm trying to factor it since there are two of the same terms here. So far this is what I have: [2x(3x+4)^1/3](3x+4)^3/3+2x]

OpenStudy (anonymous):

Okay...just give me a sec. I'm only 'half' here. You should notice 2x(3x+4)^(1/3) is a common factor.

OpenStudy (anonymous):

yeahhh I can see that. Thanks for trying to help though!

OpenStudy (anonymous):

\[2x(3x+4)^{4/3}+4x^2(3x+2)^{1/3}=2x(3x+2)^{1/3}[(3x+4)+2x]\]\[=2x(3x+2)^{1/3}(6x+4)=4x(3x+2)^{1/3}(3x+2)=4x(3x+2)^{4/3}\]

OpenStudy (anonymous):

Wait...I made a mistake in the last line...

OpenStudy (anonymous):

I'll scan in what I wrote. I never translate things onto the site properly :[

OpenStudy (anonymous):

THANK YOU SOOO MUCH! I can see how to do it now, thanks again! :)

OpenStudy (anonymous):

kk ;]

OpenStudy (anonymous):

wait, isn't 3x+2x = 5x?

OpenStudy (anonymous):

It is...damn...thanks for picking that up.

OpenStudy (anonymous):

heh, np.

OpenStudy (anonymous):

Okay...*this* one should be right. Too many distractions going on...

OpenStudy (anonymous):

I just realized something...when I wrote the problem where I have 4x^2, the original problem actually has the x^2 separated from the 4. So it actually is:\[2x(3x+4)^4/3+x^2*4(3x+4)^1/3\]

OpenStudy (anonymous):

Does this make a difference to the answer...I went ahead and multiplied x^2+4...to make it 4x^2. I hope this doesn't change all the hard work you did, I just realized this though.

OpenStudy (anonymous):

no, it doesn't matter. a * b = b * a. in other words, 7 times 9 is the same as 9 times 7 is the same as 63

OpenStudy (anonymous):

okay...thanks.

OpenStudy (anonymous):

I am getting ready to post a new thread, maybe you can help?

OpenStudy (anonymous):

maybe I can. lets see :)